A triangle is formed by the lines y=x, y=2x-1 and 3y-4x=1. Find the area of the triangle


1
Expert's answer
2021-09-02T16:38:46-0400

Given equation are "y=x,y=2x-1,3y-4x=1"

Let "y=x...(1)"

"y=2x-1\u2026(2)"

"3y-4x=1\u2026(3)"

Solving (1) and (2), we get:

"(x,y)=(1,1)"

Solving (2) and (3), we get:

"3(2x-1)-4x=1"

"\u21d22x=4" 

"\u21d2x=2"

"\u2234y=3"

"(x,y)=(2,3)"

Solving (1) and (3), we get:

"(x,y)=(-1,-1)"

So, let coordinates of points A, B, C of "\triangle ABC" be "(1,1),(2,3),(-1,-1)" respectively.

We know that area of triangle with vertices "(x_1,y_1),(x_2,y_2),(x_3,y_3)" is given as:

"\triangle =|\frac{1}{2}
\begin{vmatrix}
 x_1 & y_1 & 1 \\
 x_2 & y_2 & 1 \\
x_3 & y_3 & 1
 \end{vmatrix}|"

"\therefore" Area of "\triangle ABC=|\frac{1}{2}
\begin{vmatrix}
 1 & 1 & 1 \\
 2 & 3 & 1 \\
-1 & -1 & 1
 \end{vmatrix}|"

"=\frac{1}{2} [1(3+1)-1(2+1)+1(-2+3)]"

"=\frac{1}{2} [4-2+1]"

"=\frac{3}{2}" square units.